# LLMs hallucinate graphs too: a structural perspective

### v2

@@ 3 A LLM Hallucination Rank Based on Output Graphs
In the light of the only 5 graphs prompted to judge a LLM, and considering the 50k prompts in the Hallucination Leaderboard, we find this result to be encouraging for further study on the discriminative power of querying LLMs for structured data such as graphs.
4 Discussion
Reasoning, overfitting or hallucinating
So far, no LLM managed to consistently output the ground truth graphs. This might be considered as hallucinations, a sign that the LLMs handling of structures still needs to be improved. What is required from LLMs is yet not the overfitting and storage of input data, but rather to generalize from the provided examples. In this light, witnessing large blobs of (edgelist) text correctly returned by the LLM might be interpreted as an additional metric for overfitting.
Interestingly, the capacity to produce answers to a MCQ dataset like PiQa is interpreted as reasoning – rather than again overfitting –. In essence, reasoning could produce Les Misérables graph provided the original novel and the edge definition (chapter co-occurrence).
@@ 4 Discussion
Our approach explores graph requests as a mean to obtain more information bits by prompt, compared to the standard MCQ approach in which each request yields a number of bits capturing the number of possible answers to each question (typically 2 or 4 choices, hence 1 or 2 bits). Implicitly, this approach hence compares interrogation patterns "bit to bit", regardless of the relevance of the collected bits. A natural limitation is to overlook the precise meaning of each bit. Concretely, for instance, one bit yielded by our approach captures whether the target LLM correctly predicted a (Jean Valjean, Cosette) relation in Les Misérables, whereas one bit yielded by MedMCQA benchmarks captures whether the target LLM correctly associates "A 40-year-old man has megaloblastic anemia and early signs of neurological abnormality" to a deficit in B12 Vitamin. The difficulty with considering the semantics associated with each bit is that the notion of relevance is strongly application dependent. The relevance varies depending on the use case, be it for fingerprinting a target LLM or for evaluating its ability to provide truthful medical advice.
@@ 5 Related Work
Graphs are already used in various ways regarding the hallucination [15, 20, 19] issue, in order to assess properties or judge on the quality of outputs from LLMs.
Knowledge graphs are leveraged in a LLM-based hallucination evaluation framework [19], by prompting for text and checking the correctness of the output having a binary labeled dataset at hand. Knowledge graphs are constructed from unstructured textual data by identifying the set of entities within the text and the relationships between them, resulting in a structured representation.Work in [11] models social networks as graphs in order to simulate information spreading in order to track LLM hallucinations flowing within these networks. Nonkes et al. [16] create a graph structure that connects generations that lie closely in the embedding space of hallucinated and non-hallucinated LLM text generations. Graph Attention Networks then learn this structure and generalize it to unseen generations for categorization. Work in the complex networks community [12] is interested in extracting organizational networks (bipartite user-activity networks here) from raw text obtained from LLMs, in the context of standardized individual-level contributions across a large numbers of teams participating to a competition for instance.
To be best of our knowledge, our work is the first to propose a head-to-head comparison of a ground truth graph to a prompted output graph given by a LLM under scrutiny, allowing for new assessments.
@@ 6 Conclusion
This paper made the case for comparing ground truth graphs from the literature to the output of LLMs, prompted to output them. A first striking observation is that current LLMs are far from being reliable sources of edge lists (i.e. they do hallucinate graphs). Nevertheless, these glitches open solid comparison avenues, which we exploited to observe the significant differences of LLMs in their hallucination amplitude regarding standard graphs. Introducing relevant metrics, we have illustrated that a handful of prompts to a LLM can correlate with a method leveraging tens of thousands of queries for binary answers, in the capability to rank the less hallucinating LLMs. We believe that such a novel perspective leaves ways for future work to constitute robust and more efficient benchmarks for assessing the quality of LLMs.
@@ Appendix A Example Output From a LLM
import networkx as nx
# Define the edge list
edges = [
    (26, 33), (27, 33), (28, 33), (29, 33), (30, 33), (31, 33), (32, 33)
# Create a graph from the edge list
G = nx.Graph(edges)
# Print the graph
print("Nodes:", G.nodes())
print("Edges:", G.edges())
@@ Appendix B Statistics and Embeddings for Les Misérables and Graph Atlas 50
Statistics related to Section 2 appear in Tables 2 and 3, as well as the embeddings for these 2 graphs in Figure 5.

### v3

@@ 3 A LLM Hallucination Ranking Based on Output Graphs
In the light of the fact that we are using only five graphs to evaluate an LLM, when comparing to the use of 50k prompts in the Hallucination Leaderboard, we find this result to be encouraging for further study on the discriminative power of querying LLMs for structured data such as graphs.
4 Ability to Generate Random Graphs (RQ2)
We now investigate the ability of LLMs to generate randomness in a structured form. We choose to focus on the generation of Erdős–Rényi random graph. As a prerequisite for generating random graphs, a first natural question is whether LLMs can generate any randomness. Recent literature provides a positive answer: Hopkins et al. [17] looked at random number generation and shows that it is possible in some settings and for some LLMs to generate a rather uniform distribution on a small interval. On a different perspective, Koevering and Kleinberg [20] show that LLMs are keen to replicate human biases in random generations and at times even exacerbate them. On the other hand, Harrison Harrison [16] finds that gpt3.5 has lower repeat frequencies and adjacent number frequencies than humans.
In this work, our goal is to evaluate the randomness ability through the concept of the emergent ability and the lens of hallucination (which will connect to RQ1). To support this rationale, a simple baseline is desirable to evaluate this generative task. Consequently, we choose to focus on Erdős–Rényi random graphs, as it is arguably the simplest generative model from the literature (see e.g. Watts-Strogatz [34] or Barabasi-Albert [11] models for an increased amount of parameters).
4.1 Experimental Method
Selecting LLMs and prompts.
In RQ1, we issue a single request to LLMs to obtain the recited graph. In RQ2, as the outputs are randomized, we now require more requests to judge the quality of the generated graphs. For this reason, we choose to investigate to open weight LLMs that we run on the Jean Zay supercomputer [^3] http://www.idris.fr/jean-zay/. We list the 25 selected LLMs in Figure 5.
Prompting strategy
We explore two prompting strategies. The first strategy asks straightforwardly for a Erdős–Rényi graph realization and specifies the format of the sought output. We refer to this prompt as Direct Prompt. As prompts are known to greatly influence the LLM outputs, we also experiment with a Chain of Thought prompt CoT that generally elicits better results [37]. The precise prompts are described in appendix D.
4.2 Evaluating LLM Outputs
We now need to assess the quality of the generated outputs. We are interested in knowing if the generated graphs are representative of Erdős–Rényi graphs. To this end, we resort to a graph-level hypothesis testing strategy.
Parsing the generated outputs.
As in RQ1 setting, we leverage a regular expression based script to extract the edge list from the LLM output (see appendix D for more details). We define $\mathcal{G}(M,T,n,p)$ as the set of graphs that have been successfully extracted from the set of queries $N_{\text{queries}}$ given to the LLM $M$, with temperature $T$ [^4] The temperature of an LLM is a sampling parameter that can be set at the inference. Increasing $T$ flattens the next token probability distribution over the token vocabulary, while having $T\simeq 0$ forces the model to having a deterministic output [40]. and $(n,p)$ being the Erdős–Rényi parameters (with $n$ the number of nodes, and $p$ the probability that an edge exists between any two distinct nodes). We precise that we refer to the graphs that are “successfully" extracted, because, as in the experiments from RQ1, LLMs might not always respect the demanded format in their outputs. If this happens, our regular expression may fail to identify a graph within the LLM output. Let
[…]
be the fraction of Syntactically Correct Answers, at least with respect to our regular expression based algorithm. In our experiments, $N_{\text{queries}}$ is set to $200$.
Hypothesis Testing.
Let $H_{0}$ be: “The graph has been generated from an Erdős–Rényi process". As this (graph-level) hypothesis is not easy to test, we resort to a second (node-level) hypothesis $\tilde{H_{0}}$ whose test is more accessible. To construct $\tilde{H_{0}}$ we rely on a property of the Erdős–Rényi model: as a consequence of the graph generation process in which each edge is added independently, node degrees follow a binomial distribution. We hence define $\tilde{H_{0}}$: “Node degrees follow a binomial distribution", the rationale being that since $H_{0}\Rightarrow\tilde{H_{0}}$, rejecting $\tilde{H_{0}}$ leads to rejecting $H_{0}$.
Metric definition.
We use a $\chi^{2}$ hypothesis test, whose statistic is:
$\chi^{2}=\sum_{i=0}^{n-1}\frac{(d_{i}-n\beta_{i})^{2}}{n\beta_{i}},$ (2)
where $(d_{i})_{0\leq i\leq n-1}$ is the number of graph nodes having $i$ neighbors and $\beta_{i}=\mathbb{P}(X=i)$ (with $X\sim Bin(n-1,p)$ binomial distribution and $X$ a random variable).
[…]
Note that this test is not sufficient, since it allows to conclude that graphs failing this test are not generated from an Erdős–Rényi model with a good confidence. On the other hand, it does not guarantee that graphs passing this test come indeed from an Erdős–Rényi model. In fact, it could be possible to have a graph generation processes that would result in graphs having binomial degree distribution but that are not Erdős–Rényi .
We can define the Test Success Rate for a set of graphs $\mathcal{G}$ as the average $\chi^{2}$ test success rate:
[…]
Note that this metric needs to have a fixed number of nodes $n$ and knowledge of the distribution of $X$ (to determine $\beta_{i}$), in our case the binomial distribution for Erdős–Rényi graphs.
As a baseline for this metric, we computed the performance on graphs generated by the algorithm for the Erdős–Rényi model from the Python library NetworkX). This will be labeled as the “True Erdős–Rényi model". We note that the $\gamma$ metric is not perfect, as NetworkX achieves an average score of 97% (instead of an ideal 100%). This gap only appears for small values of $n$, which is predictable since $n$ directly determines the number of samples for each $\chi^{2}_{\text{test}}$, since there are as much samples as many node degrees. The low number of samples weakens the test, thus explaining the observed gap.
Figure 4: Test Success Rate $\gamma$ and Syntactically Correct Answer Rate $\sigma$ over LLM’s outputs on Erdős–Rényi graphs. The 7 $(n,p)$ Erdős–Rényi parameters are aggregated by mean. The experiment includes 25 LLMs and 200 graph queries per LLM-$(n,p)$ pair. Temperature is set to 1.0 and we are here using the CoT Prompt.
4.3 Experiments on the Generative Capabilities of 25 LLMs
Figure 4 represents simultaneously the syntactically correct answer rate $\sigma$ and the performance $\gamma$, given that both constitute important criteria when assessing the ability of LLMs to generate random graphs. The majority of models has a performance $\gamma$ around 45%. However the syntactically correct answers rate $\sigma$ is much more heterogeneous, with two major model clusters: one under 25% and the other above 60%. Finally five models (#4, #5, #7, #8, #9) have a really low $\sigma$, and two of them (#8, #9) fail completely the test.
We observe that there is no generalized impact of LLMs belonging to a given family of models, and no impact of the model size. However there are some model pairs of family and size that are close (#19, #20), (#2, #3), (#12, #14); though some others model pairs of family and size are far away (#21, #20), (#23, #24). We finally note that the Instruct models generally work better that non Instruct ones, as it was expected given that they have been finetuned to better respond to natural language instructions.
Figure 4 averages out the impact of different Erdős–Rényi parameters, hiding the variability they cause. The variability is highlighted in Figure 5, which shows the median performance of models on different parameters and the interquartile range, in order to underline the amplitude of the variability across models. The apparent uniformity in Figure 4 should be interpreted with caution, as there are significant variations across both parameter pairs and LLMs. Interestingly, Figure 5 shows that all models struggle to correctly produce certain Erdős–Rényi parameter pairs, with $(n=15,p=0.5)$ being the hardest.
Figure 5: Erdős–Rényi parameters $(n,p)$ influence on Test Success Rate $\gamma$, with prompt variation. The models $M$ are represented as median and interquartiles. The experiment includes 25 LLMs, 7 $(n,p)$ parameter pairs and 200 graph queries per LLM-$(n,p)$-prompt triplet. Temperature is 1.0.
We see that for the parameter pairs $(n,p\leq 1/n)$ (on the rightmost side), the CoT Prompt performs significantly better than the Direct Prompt. In contrast, for the other parameter pairs, the performance is more balanced, with the Direct Prompt performing slightly better.
Furthermore, Figure 6 represents the impact of the LLM temperature $T$ on the task. As an increase in temperature flattens the probability distribution of the next token in the output of a LLM [40], it raises the likelihood of ending up generating incoherent text. Consequently, the decreasing trend in the curve of syntactically correct answers (orange) is expected. In contrast, increasing the temperature interestingly improves the success rate $\gamma$. It should be noted that this curve is averaged over the seven tested parameter pairs $(n,p)$, and that there exist some parameter pairs for which the behavior differs.
Figure 6: Temperature influence on (degree distribution based) Test Success Rate $\gamma$ and Syntactically Correct Answer Rate $\sigma$ on Erdős–Rényi graphs queried. The $(n,p)$ Erdős–Rényi parameters dimension is averaged while the models $M$ are represented as median and interquartiles. The experiment includes 25 LLMs, 7 $(n,p)$ parameter pairs and 200 graph queries per LLM-Temperature-$(n,p)$ triplet.
Finally, Table 3 shows correlation with the Hallucination Leaderboard [2], yet a weaker correlation than what found in RQ1 (see Table 2), with a Spearman correlation of $0.2$. On the other hand, we can see that in both cases Orca-2-13b is ranked first, Qwen is ranked high, and phi-2 sits in the middle of both rankings.
[…]
#1 microsoft/Orca-2-13b microsoft/Orca-2-13b
#2 meta-llama/Llama-2-13b-chat-hf Qwen/Qwen2.5-7B-Instruct
#3 Qwen/Qwen2.5-7B-Instruct meta-llama/Llama-2-70b-chat-hf
#4 mistralai/Mixtral-8x7B-Instruct-v0.1 microsoft/phi-2
#5 microsoft/phi-2 mistralai/Mistral-7B-Instruct-v0.3
#6 mistralai/Mistral-7B-Instruct-v0.3 meta-llama/Llama-2-13b-chat-hf
#7 meta-llama/Llama-2-70b-chat-hf mistralai/Mixtral-8x7B-Instruct-v0.1
Spearman Correlation: 0.2
Table 3: Our Test Success Rate $\gamma$ vs a summarizing of a benchmark based on prompts with binary answers [2] (ranking from 2025-03-20). To be noted that $\star$ rank in [2] as been adjusted to the LLMs available in our experiment.
Discussion.
Despite the lenient nature of our metric for evaluating the randomness quality, the magnitude of its success suggests that an emergent ability is likely present. However, this interpretation must be considered with the remaining uncertainty on the number of samples in their training data, which could imply that the observed performance is merely a result of memorization.
Also, note that with LLM experiments, there are many parameters that can have significant impact on the results. They are essentially sampling parameters (temperature, top-k etc.), and the prompting strategy [24] [12]. Hence, our results offer a lower bound on the LLMs’ ability on the chosen task. Note that the models in the bottom left of Figure 4, which perform poorly, might be more successful in different tasks other than the one tested in RQ2 [24].
@@ Discussion.
Both RQ1 and RQ2 explore queries from graphs as a way to obtain more information bits per prompt, compared to the standard MCQ approach in which each request yields a number of bits capturing the number of possible answers to each question (typically 2 or 4 choices, hence 1 or 2 bits). Implicitly, RQ1 compares interrogation patterns “bit to bit", regardless of the relevance of the collected bits. A natural limitation is to overlook the precise meaning of each bit. Concretely, when one bit yielded by our approach captures whether the target LLM correctly predicted a (Jean Valjean, Cosette) relation in Les Misérables, one bit yielded by MedMCQA benchmarks captures whether the target LLM correctly associates “A 40-year-old man has megaloblastic anemia and early signs of neurological abnormality" to a deficit in B12 Vitamin. In RQ2, while a bit produced in a random graph is not very useful on its own right, when enough are collected, we are able to form a group that can be statistically evaluated. The difficulty with considering the semantics associated with each bit is that the notion of relevance is strongly application dependent. The relevance varies depending on the use case, for instance when evaluating the ability to provide truthful medical advice.
@@ 5 Related Work
Graphs are already used in various ways when it comes to studying the LLMs hallucinations [25, 32, 30], in order to evaluate the quality of their outputs.
Knowledge graphs are leveraged in a LLM-based hallucination evaluation framework [30], by prompting for text and checking the correctness of the output having a binary labeled dataset at hand. Knowledge graphs are constructed from unstructured textual data by identifying the set of entities within the text and the relationships between them, resulting in a structured representation.Work in [15] models social networks as graphs on which they simulate information spreading in order to track LLM hallucinations flowing within these networks. Nonkes et al. [26] create a graph structure that connects generations that lie closely in the embedding space of hallucinated and non-hallucinated LLM text generations. Graph Attention Networks then learn this structure and generalize it to unseen generations for categorization. Work in the complex networks community [18] is interested in extracting organizational networks (bipartite user-activity networks here) from raw text obtained from LLMs, e.g. in the context of standardized individual-level contributions across a large numbers of teams participating to a competition.
Hopkins et al. [17] investigated random number generation capabilities of LLM. They identified several factors that impact the generation quality, and show that in some conditions, some LLMs may generate rather uniform distributions on small intervals. Koevering and Kleinberg [20] show that LLMs are keen to replicate and even exacerbate human biases in random generations.
In complex network analysis, and to evaluate the quality of Erdős–Rényi graphs, Ouadah et al. [27] propose an approach based on degree variance goodness-of-fit, but their model of heterogeneous Erdős–Rényi law does not apply to our setting.
Finally, the concept of emergent ability has two definitions in literature. We chose to focus on one which is the ability to reason about novel problems zero-shot, without any direct training (Webb et al. [35]). The other being an ability that is not present in smaller-scale models but is present in large-scale models (Wei et al. [36]), that focus on the scaling complexity of the neural networks eliciting ability while the first definition focus on the ability to perform a certain task for which no example was present in the training data. Also, Schaeffer et al. [31] challenges the idea that LLMs develop novel abilities suddenly as their scale increases. The authors argue that these abilities appear emergent due to specific measurement and analysis methods rather than fundamental changes in the models. To be best of our knowledge, our work is the first to propose a head-to-head comparison of a ground truth graph to a prompted output graph, or a statistical evaluation of the generative capabilities of LLMs. This was performed under the scope of hallucinations, and allows for novel assessments.
@@ 6 Conclusion
The scope of this article is to compare the ability of LLMs to (1) recite ground truth graphs from the literature and (2) generate Erdős–Rényi random graphs. A first striking observation is that current LLMs are far from being reliable sources (i.e. they do hallucinate graphs) in both tasks. Nevertheless, these glitches open solid comparison avenues, which we exploited to observe the significant differences of LLMs in their hallucination amplitude regarding standard graphs. We introduced two relevant metrics for the tasks and we showed that a handful of prompts can correlate with a method leveraging tens of thousands of queries for binary answers, with the goal to rank the LLMs that hallucinating the least. For what concerns the ability to generate random graphs, the generative quality we measured is high enough to raise intriguing questions and motivate deeper exploration, e.g. starting with an extension to other generative models. We believe that such a novel perspective leaves ways for future work to better evaluate LLMs on both hallucination and randomness generative abilities.
@@ Appendix B Statistics and Embeddings for Les Misérables and Graph Atlas 50
Statistics related to Section 2 appear in Tables 4 and 5, as well as the embeddings for these 2 graphs in Figure 7.
@@ Appendix C Spectral Distances from Ground Truth Graphs
Appendix D Experiment settings
D.1 Prompts for Erdős–Rényi generations
Here are the two prompts used to generate Erdős–Rényi graphs with LLMs. The $n$ and $p$ parameters are blue-marked to signify they vary (since we adjust the tested parameters). As an example, we used $n={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}15}$ and $p={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.3}$.
Direct Prompt:
Give a realization of an Erdős–Rényi graph of parameters n=15, p=0.3.
Use only this format: edges are enclosed in parentheses, integers are separated by a comma, and edges are separated by commas. Do not include any additional explanations or text, only the edges of the graph in the exact specified format.
CoT Prompt (Chain of Thought):
Give a realization of a Erdős–Rényi graph of parameters n=15, p=0.3.
Think step by step. At the end of your reasoning, please say "final answer: " and use this format for the final answer: edges are enclosed in parentheses, integers are separated by a comma, and edges are separated by commas.
D.2 Sampling parameters
The temperature used is always mentionned on the figures (varies from 0.2 to 2.0 with a .2 step). The top-k parameter is set to $50$ and the top-p to $1$ which are default values of the Transformers library (https://huggingface.co/docs/transformers/en/main_classes/text_generation).
D.3 Regex based script
For the Direct Prompt, the script returns the first matching with an edge list of format like (1,3) (2,8) (7,2). For the CoT Prompt the script first searches for the occurence of final answer: (or similar) and then search for an edge list like above.
D.4 List of Abbreviations
  • LLM: Large Language Model
  • MCQ: Multiple Choice Question
  • KC: Karate Club
  • LM: Les Misérables
  • GAD: Graph Atlas Distance
  • CoT: Chain of Thought
  • GNN: Graph Neural Network